OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3)

OpenStudy (precal):

your slope for a perpendicular line is opposite reciprocal a and -1/a your slope is 2 the opposite reciprocal is -1/2 now use (6,3) to find your equation

OpenStudy (anonymous):

umm and how do you do that?

OpenStudy (anonymous):

Slope of given line is 2. So, slope of perpendicular line will be -1/2. So, the new line is: y = -x/2 + b. Find b using point given 6,3 for x and y. Then, you have your line.

OpenStudy (anonymous):

how did yu get that

OpenStudy (anonymous):

3 = -6/2 + b 3 = -3 + b b = 6. So, line is: y = -x/2 + 6

OpenStudy (anonymous):

Perpendicular lines have slopes that multiply to -1. So, if slope m1 = 2 and its product with m2 = -1, you have 2 * m2 = -1 => m2 = -1/2

OpenStudy (anonymous):

oK thanks